OpenStudy (anonymous):

The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting. A) 19 B) 33 C) 73 D) 84

OpenStudy (anonymous):

@smilekbug @Nnesha @uri @amistre64 @AdmiralAngel25

OpenStudy (anonymous):

Use the first formula to find out how many pizzas are needed p(9) = 9(2) + 1 = 19 Use this formula to find out how much money is needed r(19) = 4(19) - 3 = 73 Answer is C

OpenStudy (anonymous):

Thank you! Can you help me with this one to? @AdmiralAngel25 Given f(x) = the quantity of 3x minus 1, divided by 2, solve for f-1(4). 1 3 5 7

OpenStudy (anonymous):

@amistre64 @Compassionate

OpenStudy (anonymous):

One second this is tricker for me

OpenStudy (anonymous):

Something tells me its either C or D, probably C though...sorry but can't figure it out.

OpenStudy (anonymous):

its okay, thanks anyways, sorry i have one more if thats okay

OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)). @AdmiralAngel25 5 7 9 10

OpenStudy (anonymous):

f(3) = 2(3)+1 6+1 7 g(7) = \[\frac{ 3(7)-1}{ 2 }\] g(x) = 10

OpenStudy (anonymous):

ohhh ok, i understand now, thanks again!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Do you know hoe to find average rate of change for a function? @AdmiralAngel25